# A Generalized Sylvester-Fermat-Torricelli problem with application in disaster relief operations by UAVs

### v1

@@ 4 Existence and Uniqueness of Optimal Solutions in the Generalized SFT Problem and Its Weighted Variant
In the following Theorem, we provide the condition under which the existance of the solution will be proved.
Theorem 4.1 (Existence of solutions to the generalized SFT problem).
In the setting of the generalized SFT problem (2.7), an optimal solution exists if at least one of the sets $\Omega_{0}$ or $\Omega_{i}$ (for some $i\in I_{k}$, $k=1,\dots,n$) is bounded.
@@
The sublevel set for $\lambda\geq 0$ and the function $S(x)$ defined in (2.8) is given by
$\mathcal{L}_{\lambda}(S):=\{x\in\Omega_{0}\mid S(x)\leq\lambda\},$
which implies that $\sum_{i\in I_{k}}\rho_{F_{i}}^{\Omega_{i}}(x)\leq\lambda$ for all $k=1,2,\dots,n$, whenever $x\in\Omega_{0}$. Therefore, for the set
$A:=\left\{x\in\Omega_{0}\mid\rho_{F_{i}}^{\Omega_{i}}(x)\leq\lambda\ \text{for all}\ i\in I_{k},\ k=1,2,\dots,n\right\},$
we have $\mathcal{L}_{\lambda}(S)\subseteq A$. For a closed set $\Omega_{i}$, the condition $\rho_{F_{i}}^{\Omega_{i}}(x)\leq\lambda$ implies that $x\in\Omega_{i}+\lambda F_{i}$. Therefore, by the definition of $A$, it follows that
@@
Hence, if $\Omega_{i}$ is bounded for some $i\in I_{k}$ with $k=1,2,\dots,n$, then the set $A$ is bounded. Since $A$ is bounded, $\mathcal{L}_{\lambda}(S)$ is also bounded. Then, by Corollary 7.10 in [22], the existence of an optimal solution is guaranteed.
In the following theorem, under certain assumptions, we establish the uniqueness of the optimal solution in the generalized SFT problem.
Theorem 4.2 (Uniqueness of solutions to the generalized SFT problem).
In the setting of the generalized SFT problem (2.7), a unique solution exists under the following conditions:
  1. i)
    The sets $F_{i}$ and $\Omega_{i}$ are strictly convex for $i\in I_{k}$ with $k=1,2,\dots,n$.
  2. ii)
    For each index set $I_{k}$ with $k\in\{1,2,\dots,n\}$, among the sets $\Omega_{i}$ with $i\in I_{k}$, there exists $\Omega_{j}$ such that for all distinct points $x,y\in\Omega_{0}$, the line $L(x,y)$ does not intersect $\Omega_{j}$; that is, $L(x,y)\cap\Omega_{j}=\emptyset$.
  3. iii)
    At least one of the set $\Omega_{0}$ or the sets $\Omega_{i}$ for $i\in I_{k}$ with $k=1,\dots,n$ is bounded.
@@
Based on $(iii)$ and Theorem 4.1, we know that there is an optimal solution. Based on $(i)$, $(ii)$, and Proposition 8.18, each $\sum_{i\in I_{k}}\rho_{F_{i}}^{\Omega_{i}}(x)$ is strictly convex for $k=1,2,\dots,n$. It follows from Lemma 2.2 that $S(x)$ is strictly convex, thereby completing the proof.
Motivated by prior works that studied the weighted generalized Fermat–Torricelli and Sylvester problems [25, 35, 36], we conclude this section by presenting and discussing the weighted formulation of the generalized SFT problem, given by (considering the setting of the problem in (2.7)):
$\min\ \tilde{S}(x)\quad\text{subject to}\quad x\in\Omega_{0},$ (4.2)
where, for weights $\omega_{i}>0$ for all $i\in I_{k}$ and $k=1,2,\dots,n$, the objective function is defined as:
$\tilde{S}(x):=\max\{\sum_{i\in I_{k}}\omega_{i}\rho_{F_{i}}^{\Omega_{i}}(x)\ |\ k=1,2,\dots,n\}.$ (4.3)
It is worth noting that the original formulation of the generalized SFT problem in (2.7) inherently includes its weighted counterpart. By applying Proposition 3.7, we obtain
@@
which allows us to express the weighted version in the same form as the original problem:
$\min\ S(x)\quad\text{subject to}\quad x\in\Omega_{0},$ (4.4)
where
$S(x):=\max\{\sum_{i\in I_{k}}\rho_{\tilde{F}_{i}}^{\Omega_{i}}(x)\ |\ k=1,2,\dots,n\},$ (4.5)
and $\tilde{F}_{i}:=F_{i}/\omega_{i}$.
5 An Extended Formulation of the Generalized SFT Problem
@@
In this section, we give an extended version of the generalized SFT problem that incorporates both the set-based Minkowski gauge and the MSMG functions. In addition to the setting of the generalized SFT problem (2.7) consider the following settings. Let $\bar{m}\in\mathbb{N}$ and the index sets $J_{k}\neq\emptyset$ for $k=1,2,\dots,n$ form a partition of the set $J:=\{1,2,\dots,\bar{m}\}$, such that $\bigcup_{k=1}^{n}J_{k}=J$ and $J_{k}\cap J_{l}=\emptyset$ for all $k\neq l$ with $k,l\in\{1,2,\dots,n\}$. For $j\in J_{k}$ with $k=1,2,\dots n$, let $\bar{F}_{j}\subset\mathbb{R}^{q}$ be a compact, convex set containing the origin in its interior point (i.e., $0\in\operatorname{int}\bar{F}_{j}$). Let $\bar{F}_{j}\subset\mathbb{R}^{q}$ be a compact, convex set containing the origin in its interior point (i.e., $0\in\operatorname{int}\bar{F}_{j}$). Also, let $\Theta_{j}$ be nonempty compact convex subsets of $\mathbb{R}^{q}$. Then, the following problem is formulated as an extension of the generalized SFT problem:
$\min\hat{S}(x)\ \ \ s.t.\ \ \ x\in\Omega_{0},$ (5.1)
where
$\hat{S}(x):=\max\{\sum_{i\in I_{k}}\rho_{F_{i}}^{\Omega_{i}}(x)+\sum_{j\in J_{k}}\bar{\rho}_{\bar{F}_{j}}^{\Theta_{j}}(x)\ |\ k=1,2,...,n\}.$ (5.2)
@@
The application of the extended generalized SFT problem (5.1) is demonstrated in a real-world scenario in Section 6.
@@
In the case $k=1$, $|I_{1}|=m$, and $|J_{1}|=\bar{m}$ in problem (5.1), the extended version of the generalized Fermat-Torricelli problem is given by:
$\min\hat{T}(x)\ \ \ s.t.\ \ \ x\in\Omega_{0},$ (5.3)
where
$\hat{T}(x):=\sum_{i=1}^{m}\rho_{F_{i}}^{\Omega_{i}}(x)+\sum_{j=1}^{\bar{m}}\bar{\rho}_{\bar{F}_{j}}^{\Theta_{j}}(x).$ (5.4)
For the generalized Fermat-Torricelli problem in (2.11), a similar problem was studied by [12]. However, the problem proposed in (5.3) represents a new extension of the Fermat-Torricelli problem, and its application in disaster relief operations using UAVs is discussed in Section 6.3.
Finally, in the setting of the problem (5.1), we introduce the extended formulation of the generalized Sylvester problem:
$\min\hat{Y}(x)\quad\text{s.t.}\quad x\in\Omega_{0},$ (5.5)
where
$\hat{Y}(x)=\{\mathcal{Y}_{1}(x),\mathcal{Y}_{2}(x)\},$ (5.6)
with
@@
$\mathcal{Y}_{1}(x)=\max_{i=1,2,\dots,m}\rho_{F_{i}}^{\Omega_{i}}(x),\quad\mathcal{Y}_{2}(x)=\max_{j=1,2,\dots,\bar{m}}\bar{\rho}_{F_{j}}^{\Theta_{j}}(x).$ (5.7)
It should be noted that both problems proposed in (2.9) and (5.5) constitute novel generalizations of the Sylvester problem. Their application in disaster relief operations using UAVs is discussed in Sections 6.2 and 6.3.
The convexity of the functions $\hat{S}(x)$, $\hat{T}(x)$, and $\hat{Y}(x)$—defined in (5.2), (5.4), and (5.6), respectively—follows directly from Proposition 3.2, Proposition 2.1, and Theorem 3.9. Furthermore, their subdifferentials and subgradients can be derived using Propositions 3.4, 3.5, 3.6, together with Theorems 3.9, 2.3, and 2.4.
6 Application of the Generalized SFT Problem to Optimizing Facility Location Decisions in Disaster Relief Operations
@@
After an earthquake, some people may be trapped under debris but still alive. If they have access to their cell phones, one of the most effective ways they can help themselves is by sending their location to family members or posting it on social media. This can greatly increase their chances of being found and rescued. However, earthquakes often damage or destroy communication systems, such as cell towers and internet cables. As a result, even if someone has a working phone, it may not be able to send messages or connect to the internet.
@@
In such cases, UAVs can play an important role. By flying over the damaged area, UAVs can provide wireless connections or DTNs, allowing trapped individuals to send their location information (see Fig. 1). This support can make rescue operations faster and more effective, and help save more lives. Despite their advantages, UAVs also present several challenges. They are limited by short battery life and are vulnerable to environmental conditions, especially high winds. Moreover, rescue teams often deploy a heterogeneous fleet, consisting of UAVs with varying speeds, ranges, and operational capabilities, which adds complexity to coordination and mission planning.
@@
In this section, we examine several scenarios in which a truck serves as a mobile station for deploying and recharging UAVs. The primary objective is to determine the optimal location of the truck (see Fig. 1) to minimize the transition time—the period when the UAV is traveling between the service area and the mobile station, resulting in a temporary loss of network service. This problem can be formulated as a single-facility location problem, which has been extensively studied in the literature using the $\ell_{1}$ and $\ell_{2}$ norms. However, to model a more realistic scenario, we incorporate the effect of wind, which prevents us from using these common norms directly. Using the $\ell_{2}$ norm, for example, would ignore the wind’s influence, leading to a simplification that can result in significant time loss. To minimize transition time, we use the set-based Minkowski gauge function (2.6) that captures the effects of wind, which are inherent in real-world conditions.
@@
In the first scenario, a rescue team utilizes a single UAV, and we show how the generalized Fermat-Torricelli problem can be applied to minimize transition time. In the second scenario, the rescue team operates a heterogeneous fleet of UAVs, and we illustrate how the generalized Sylvester problem can effectively address the complexities that arise. Finally, we consider the generalized SFT problem in a scenario where the heterogeneous fleet of UAVs is first deployed to the affected area and then returns to the truck for recharging and data transfer. This setup highlights the operational challenges and the importance of strategic truck placement in dynamic rescue missions.
@@
6.1 Application of the Generalized Fermat-Torricelli Poblem
Owing to constraints such as limited budget, it is often the case that a rescue team operates with only a single UAV. In this section, we examine an operation in which the rescue team employs a UAV to provide DTN support for three predetermined areas. First, the UAV is dispatched from the mobile station (e.g., a truck) and covers the first area. After collecting data, the UAV returns to the station to recharge and transfer the collected data to the rescue team. This pattern is then repeated for the remaining two areas. To determine the optimal location for the truck—minimizing transition time under windy conditions—we apply the generalized Fermat-Torricelli problem, a variation of the generalized SFT problem. We present a step-by-step solution to illustrate how the generalized SFT problem is applied in this context.
To formulate the stated challenge (corresponding to Case 1 in Table 1) as a generalized Fermat-Torricelli problem (2.11), we begin by specifying the reference and dynamic sets. The reference sets are:

### v2

@@ 4 Well-Posedness of the Generalized SFT Model
The first theorem establishes the existence of optimal solutions to the generalized SFT problem (2.6) under a simple assumption.
Theorem 4.1.
In the setting of the generalized SFT problem (2.6), an optimal solution exists if at least one of the sets $\Omega_{0}$ and $\Omega_{i}$ as $i\in I_{k}$, $k=1,\dots,n$, is bounded.
@@
For any $\lambda\geq 0$, define the sublevel set associated with the function $S(x)$ from (2.7) relative to the constraint set $\Omega_{0}$ in (2.6) by
$\mathcal{L}_{\lambda}(S):=\big\{x\in\Omega_{0}\;\big|\;S(x)\leq\lambda\big\}.$
This implies that $\sum_{i\in I_{k}}\rho_{F_{i}}^{\Omega_{i}}(x)\leq\lambda$ as $k=1,\dots,n$ whenever $x\in\Omega_{0}$. Therefore, we have $\mathcal{L}_{\lambda}(S)\subset A$ with the set $A$ given by
$A:=\big\{x\in\Omega_{0}\;\big|\;\rho_{F_{i}}^{\Omega_{i}}(x)\leq\lambda\ \text{for all}\ i\in I_{k},\ k=1,\dots,n\big\},$
Since the sets $\Omega_{i}$ are closed, it follows from $\rho_{F_{i}}^{\Omega_{i}}(x)\leq\lambda$ that $x\in\Omega_{i}+\lambda F_{i}$. Thus
@@
This tells us that the boundedness of one of the sets $\Omega_{i}$ for some $i\in I_{k}$ with $k=1,\dots,n$ yields the boundedness of the set $A$, and hence this property of the sublevel set $\mathcal{L}_{\lambda}(S)$. The existence of optimal solutions to (2.6) follows now from [mordukhovich2023easy, Corollary 7.10] due to the classical Weierstrass existence theorem. ∎
Before deriving the uniqueness result for (2.6), we present the following lemma of its independent interest, which is used in the proof of the uniqueness theorem below.
Lemma 4.2.
In addition to the standing assumptions in the formulation of the generalized SFT problem (2.6), suppose that the sets $F_{i}$ and $\Omega_{i}$, $i=1,\dots,m$, are strictly convex. If furthermore for any line $L(x,y)$ from (4.1) with $x,y\in\Omega_{0}$ and $x\neq y$, there exists $j\in\{1,\dots,m\}$ such that $L(x,y)\cap\Omega_{j}=\emptyset$, then the function $T(x)$ defined in (2.11) is strictly convex on the set $\Omega_{0}$.
@@
It follows from Propositions 2.1, 3.2 that $T(x)$ is convex. Suppose on the contrary that $T(x)$ is not strictly convex. Then there exist $x,y\in\Omega_{0}$ with $x\neq y$ and $\lambda\in(0,1)$ such that
$T(\lambda x+(1-\lambda)y)=\lambda T(x)+(1-\lambda)T(y).$
Since each $\rho_{F_{i}}^{\Omega_{i}}$ is convex, we get by definition that
$\rho_{F_{i}}^{\Omega_{i}}(\lambda x+(1-\lambda)y)\leq\lambda\rho_{F_{i}}^{\Omega_{i}}(x)+(1-\lambda)\rho_{F_{i}}^{\Omega_{i}}(y)\;\text{ whenever }\ i=1,\dots,n.$
As assumed, there exists $j$ with $L(x,y)\cap\Omega_{j}=\emptyset$. Pick $\omega\in\Pi_{F_{j}}(x;\Omega_{j})$, $v\in\Pi_{F_{j}}(y;\Omega_{j})$ and then write $\rho_{F_{j}}^{\Omega_{j}}(x)=\rho_{F_{j}}(x-\omega)$ and $\rho_{F_{j}}^{\Omega_{j}}(y)=\rho_{F_{j}}(y-v)$. This gives us the relationships
$\begin{split}\lambda\rho_{F_{j}}(x-\omega)+(1-\lambda)\rho_{F_{j}}(y-v)&=\lambda\rho_{F_{j}}^{\Omega_{j}}(x)+(1-\lambda)\rho_{F_{j}}^{\Omega_{j}}(y)\\ &=\rho_{F_{j}}^{\Omega_{j}}(\lambda x+(1-\lambda)y)\\ &\leq\rho_{F_{j}}(\lambda x+(1-\lambda)y-(\lambda\omega+(1-\lambda)v))\\ &\leq\lambda\rho_{F_{j}}(x-\omega)+(1-\lambda)\rho_{F_{j}}(y-v),\end{split}$
which imply therefore that $\lambda\omega+(1-\lambda)v\in\Pi_{F_{j}}(\lambda x+(1-\lambda)y;\Omega_{j})$. We know by Proposition 8.2 that $\omega\in\textnormal{bd}(\Omega_{j})$, $v\in\textnormal{bd}(\Omega_{j})$, and $\lambda\omega+(1-\lambda)v\in\textnormal{bd}(\Omega_{j})$. Since the set $\Omega_{j}$ is strictly convex, it follows that $\omega=v$. Furthermore, the function $\rho_{F}$ is positively homogeneous, and hence
By $x,y\notin\Omega_{j}$ we get $x-\omega,y-\omega\neq 0$. Moreover, Proposition 8.3 tells us that $\alpha>0$ with $\lambda(x-\omega)=\alpha(1-\lambda)(y-\omega)$. This brings us to the representation
$x-\omega=\beta(y-\omega)\ \ \textnormal{for}\ \ \beta=\frac{\alpha(1-\lambda)}{\lambda}\neq 1,$
[…]
Finally, we are ready to establish the uniqueness of optimal solutions to the generalized SFT problem (2.6) under appropriate assumptions.
Theorem 4.3.
In addition to the standing assumptions on the data of (2.6), suppose that:
(i) The sets $F_{i}$ and $\Omega_{i}$ are strictly convex for $i\in I_{k}$ with $k=1,\dots,n$.
(ii) For each index set $I_{k}$ with $k\in\{1,\dots,n\}$, there exists a set $\Omega_{j}$ among $\Omega_{i}$ with $i\in I_{k}$ such that $L(x,y)\cap\Omega_{j}=\emptyset$ whenever $x,y\in\Omega_{0}$ with $x\neq y$.
(iii) At least one of the sets $\Omega_{0}$ and $\Omega_{i}$, $i\in I_{k}$, $k=1,\dots,n$, is bounded.
Then problem (2.6) admits a unique optimal solution.
Proof.
The existence of solutions to (2.6) is proved in Theorem 4.1. To verify the uniqueness, we get from Lemma 4.2 that the sum $\sum_{i\in I_{k}}\rho_{F_{i}}^{\Omega_{i}}(x)$ is strictly convex for each $k=1,\dots,n$. Furthermore, Proposition 2.2 tells us that the function $S(x)$ is strictly convex as well, which therefore ensures the claimed solution uniqueness. ∎
5 Weighted and Extended Formulations of the Generalized SFT Problem
Motivated by prior works that studied the weighted generalized Fermat–Torricelli and Sylvester problems [mordukhovich2012applications, nickel2003approach, plastria2009asymmetric, PLASTRIA201698, PELEGRIN1985327], we conclude this section by presenting and discussing the weighted formulation of the generalized SFT problem given, in the setting of (2.6), by
$\min\ \widetilde{S}(x)\quad\text{subject to}\quad x\in\Omega_{0},$
where the objective function is defined, with the weights $\omega_{i}>0$ for $i\in I_{k}$ and $k=1,\dots,n$, as
$\widetilde{S}(x):=\max\Big\{\sum_{i\in I_{k}}\omega_{i}\rho_{F_{i}}^{\Omega_{i}}(x)\ \Big|\ k=1,\dots,n\Big\}.$ (5.1)
It is worth noting that the original formulation of the generalized SFT problem in (2.6) inherently includes its weighted counterpart. Indeed, it follows from Proposition 3.7 that
@@ 5 Weighted and Extended Formulations of the Generalized SFT Problem
which allows us to express the weighted version in the same form as the original problem (2.6) where the (modified) cost function is given by
$S(x):=\max\Big\{\sum_{i\in I_{k}}\rho_{\widetilde{F}_{i}}^{\Omega_{i}}(x)\ \Big|\ k=1,\dots,n\Big\}\;\mbox{ with }\;\widetilde{F}_{i}:=F_{i}/\omega_{i}.$
This facilitates the investigation of well-posedness for the weighted generalized SFT problem using Section 4.
@@
Furthermore, we can explore the following extended version of the generalized SFT problem that incorporates both the set-based Minkowski gauge and the MSMG functions. Let $\bar{m}\in\mathbb{N}$ and the index sets $J_{k}\neq\emptyset$ for $k=1,\dots,n$ form a partition of $J:=\{1,\dots,\bar{m}\}$ such that
$\bigcup_{k=1}^{n}J_{k}=J\;\mbox{ and }\;J_{k}\cap J_{l}=\emptyset\;\mbox{ whenever }\;k\neq l\;\mbox{ with }\,k,l\in\{1,\dots,n\}.$
For $j\in J_{k}$ with $k=1,\dots n$, let $\bar{F}_{j}\subset\mathbb{R}^{q}$ be a compact and convex set with $0\in\operatorname{int}\,\bar{F}_{j}$, and let $\Theta_{j}$ be nonempty, compact, and convex subsets of $\mathbb{R}^{q}$. The extension of (2.6) is formulated by
$\min\;\widehat{S}(x)\;\mbox{ subjet to }\;x\in\Omega_{0},$ (5.2)
where the extended cost function is defined by
$\widehat{S}(x):=\max\Big\{\sum_{i\in I_{k}}\rho_{F_{i}}^{\Omega_{i}}(x)+\sum_{j\in J_{k}}\bar{\rho}_{\bar{F}_{j}}^{\Theta_{j}}(x)\ \Big|\ k=1,...,n\Big\}.$ (5.3)
@@
Applications of the generalized SFT problem in the extended form (5.2) are demonstrated in a real-world scenario in Section 6.
@@
In the case where $k=1$, $|I_{1}|=m$, and $|J_{1}|=\bar{m}$ in problem (5.2), the extended version of the generalized Fermat-Torricelli problem is represented by
$\min\widehat{T}(x)\;\mbox{ subject to }\;x\in\Omega_{0}$ (5.4)
with the cost function given in the summation form
$\widehat{T}(x):=\sum_{i=1}^{m}\rho_{F_{i}}^{\Omega_{i}}(x)+\sum_{j=1}^{\bar{m}}\overline{\rho}_{\overline{F}_{j}}^{\Theta_{j}}(x).$ (5.5)
Applications of (5.4), (5.5) to disaster relief operations using UAVs are discussed in Section 6.3.
Finally in this section, we formulate an extended version of the generalized Sylvester problem given by
$\min\widehat{Y}(x)\;\mbox{ subject to }\;x\in\Omega_{0},$ (5.6)
where both components of the cost functions are represented in the maximun function form
@@
$\widehat{Y}(x)=\big\{\mathcal{Y}_{1}(x),\mathcal{Y}_{2}(x)\big\}\;\mbox{ with }\;\mathcal{Y}_{1}(x):=\max_{i=1,\dots,m}\rho_{F_{i}}^{\Omega_{i}}(x),\quad\mathcal{Y}_{2}(x):=\max_{j=1,\dots,\bar{m}}\overline{\rho}_{F_{j}}^{\Theta_{j}}(x).$ (5.7)
Both versions of the generalized Sylvester problem proposed in (2.8) and (5.6) constitute novel models whose applications to UAV-based disaster relief operations are discussed in Sections 6.2 and 6.3. Observe that the convexity of the functions $\widehat{S}(x)$, $\widehat{T}(x)$, and $\widehat{Y}(x)$, defined in (5.3), (5.5), and (5.7), respectively, follows directly from Propositions 3.2, 2.1 and Theorem 3.9. Furthermore, their subgradient evaluations can be obtained by using Theorems 3.4, 3.5 and Proposition 3.6 combined with Theorems 3.9, 2.3, and 2.4.
6 Applications of the New Location Science Models to Disaster Relief Operations
@@
After an earthquake, some people may be trapped under debris but still alive. If they have access to their cell phones, one of the most effective ways they can help themselves is by sending their location to family members or posting it on social media (see, e.g., [toraman2023tweets] for a real case). This can greatly increase their chances of being found and rescued. However, earthquakes often damage or destroy communication systems such as cell towers and internet cables. As a result, even if someone has a working phone, it may not be possible to send messages or connect to the Internet.
@@
In such cases, UAVs may play a crucial role. By flying over the damaged area, UAVs can provide wireless connections or DTNs (delay-tolerant networks) while allowing trapped individuals to send their location information (see Fig. 2). This support is able make rescue operations faster and more effective, and thus help saving more lives. Despite their advantages, UAVs also present several challenges; in particular, they are limited by short battery life and are vulnerable to environmental conditions, especially high winds. Moreover, rescue teams often deploy a heterogeneous fleet consisting of UAVs with varying speeds, ranges, and operational capabilities adding complexity to coordination and mission planning.
@@
In this section, we examine several scenarios in which a truck serves as a mobile station for deploying and recharging UAVs. The primary objective is to determine the optimal location of the truck (see Fig. 2) to minimize the transition time, i.e., the period when the UAV is traveling between the service area and the mobile station, resulting in a temporary loss of network service. Such models can be formulated via single-facility location problems, which have been extensively studied in the literature by using the $\ell_{1}$ and $\ell_{2}$ norms. However, to model a more realistic scenario, we incorporate the effect of wind that prevents us from employing these common norms directly. Using the $\ell_{2}$ norm, for example, would ignore the wind’s influence, leading us to a simplification that can result in significant time loss.
@@
To minimize transition time, we use the set-based Minkowski gauge function (2.5) that captures the effects of wind, which are inherent in real-world conditions. In the first scenario, a rescue team utilizes a single UAV, and we show how the generalized Fermat-Torricelli problem can be applied to minimize transition time. In the second scenario, the rescue team operates a heterogeneous fleet of UAVs, and we illustrate how the generalized Sylvester problem can effectively address the complexities that arise. Finally, we consider the generalized SFT problem in a scenario where the heterogeneous fleet of UAVs is first deployed to the affected area and then returns to the truck for recharging and data transfer. This setup highlights the operational challenges and the importance of strategic truck placement in dynamic rescue missions.
@@
6.1 Applications of the Generalized Fermat-Torricelli Problem
Owing to constraints such as limited budget, it is often the case that a rescue team operates with only a single UAV. In this section, we examine an operation in which the rescue team employs a UAV to provide DTN support for three predetermined areas. To begin with, the UAV is dispatched from the mobile station (e.g., a truck) and covers the first area. After collecting data, the UAV returns to the station for recharging and transferring the collected data to the rescue team. This pattern is then repeated for the remaining two areas. To determine the optimal location of the truck in order to minimize transition time under windy conditions, we apply the above generalized Fermat-Torricelli model. We present a step-by-step solution to illustrate how our model is implemented in this context.
To formulate the aforementioned challenge (corresponding to Case 1 in Table 1) as a generalized Fermat-Torricelli problem (2.10), we begin by specifying the reference and dynamic sets. The reference sets are given as follows: